# The Impact of Copyrighted Material on Large Language Models: A Norwegian Perspective

### v3

@@ 1 Introduction
However, the vast quantities of data required for training these models often include copyrighted materials, presenting novel challenges related to intellectual property rights and compensation. Additionally, prior research has highlighted significant concerns about dataset composition and quality in large-scale web-crawled datasets, emphasizing the need for more responsible data curation practices (Kreutzer et al., 2022; Artetxe et al., 2022; Penedo et al., 2024). Together, these challenges have led to numerous lawsuits across jurisdictions, fundamentally questioning the legitimacy of training models on copyrighted data without explicit permissions from content creators (Panettieri, 2024; Madigan, 2024; Weisenberger et al., 2024) [^1] See Gervais et al. (2024) for an in-depth introduction on how LLMs are being interpreted in the legal domain..
The first wave of lawsuits emerged shortly after the public release of advanced generative AI models (see Appendix A). Content creators, including authors, visual artists, and musicians, began to express concerns about the unauthorized use of their work in training datasets. Multiple class-action lawsuits were filed in the United States, accusing prominent AI companies such as OpenAI and Meta Platforms of infringing on copyright laws by using copyrighted materials without obtaining explicit permissions. The authors argued that the unauthorized use of their works without any form of compensation or recognition undermines their intellectual property rights and jeopardizes their ability to earn a living from their creative endeavors. In Europe, a coalition of news publishers has taken legal action against Google and Meta Platforms, arguing that the use of journalistic content in training models without fair remuneration constitutes a breach of copyright and undermines the sustainability of high-quality journalism. Likewise, Norwegian rights-holder organizations representing publishing houses across the country, contacted the government in late 2023 expressing their concerns over the use of their material in training generative language models and demanding some sort of compensation were their contents to be used in the training of generative language models. As a result, the Ministry of Culture and Equality (Kultur- og likestillingsdepartemente) instructed the National Library to create a data-driven report they could use in order to make informed decisions in the elaboration of a compensation scheme for the authors. Led by the National Library of Norway, a consortium was formed together with the University of Oslo and the Norwegian University of Science and Technology under the umbrella of the so-called Mímir Project [^2] A name chosen after a figure in Norse mythology renowned for his knowledge and wisdom..
@@ 3.1 Core Datasets
We followed the recipe of the Norwegian Colossal Corpus (NCC) by Kummervold et al. (2022), adapting and updating it with new up-to-date contents, re-OCRing some materials, enriching their metadata, and ensuring uniform format and functionality across datasets. The preparation involved cleaning, deduplication, metadata tagging, and language balancing to maintain consistent representation of Norwegian, preventing other languages from overshadowing it. The corpus was divided into two main datasets: a base dataset excluding publisher-controlled copyright-protected books and newspapers [^3] Except for newspapers that fall under the Language Technology Use (Språkteknologiformål), as they were already included in other datasets such as NCC., and an extended dataset that included all collected texts, thus including all of base (see Table 1).
@@
We decided to include texts from other Scandinavian (Swedish, Danish, and Icelandic) and English sources to boost the performance of the resulting language models via cross-lingual transfer (Conneau et al., 2020b; Xue et al., 2021). To ensure that languages other than Norwegian, and primarily coming via Internet crawling, were balanced, we adapted the perplexity-based sampling strategy from De la Rosa et al. (2022) to maintain a high quality in the selected data. Instead of sampling a fixed number of documents, parameters for a Gaussian curve were calculated from 500,000-1M random documents per source, utilizing Wikipedia-based Kneser-Ney language models from Wenzek et al. (2019) and Conneau et al. (2020a) [^4] Built with KenLM Heafield (2011).. We also modified the perplexity calculation to account for normalized text. These parameters then guided dataset sub-sampling to target ratios per language, reducing foreign language content while maintaining quality (see Appendix B).
@@ 3.2 Domain Specific Subsets
The publisher-controlled copyright-protected materials present in the extended dataset were further divided into groups attending to different criteria. These subsets were carefully designed to test the effect of adding them to the training sets for LLMs. We split the books into fiction vs nonfiction, and original works in Norwegian vs translations. While most books in the collection had metadata information regarding the original language in which a work was written in, genre labels were more scarce. To overcome this limitation, we built a Doc2Vec model [^5] https://huggingface.co/mimir-project/literary-form-classifier Le and Mikolov (2014) that classified fiction vs nonfiction with 98% accuracy and used it to annotate books for which this information was missing. As shown in Table 2, we then built domain specific subsets to investigate 1) the effect of books vs newspapers vs books + newspapers, 2) the effect of factuality by adding only fiction words, only nonfiction works, and nonfiction works + newspapers, and 3) the effect of adding content written originally in Norwegian, such as original books or original books + newspapers, vs translated books.
@@ 3.3 Instruction-tuning Datasets
To align the models more closely with human objectives and assess whether instruction tuning with limited high-quality data can enhance the performance of our pre-trained models across various tasks, we built upon prior work and collected nearly 5,000 instructions annotated by research assistants [^6] https://huggingface.co/datasets/mimir-project/mimir-instruction. The instructions were formatted as (instruction, input, output) triplets, where instruction refers to the directive provided by humans for the model, input is an optional field containing task-related information, and output denotes the desired response that follows the given instruction.
@@ 4 Model Training
    To measure the overall impact of publisher-controlled copyrighted corpora and its effect in realistic scenarios, we conducted pre-training on the base and extended datasets, both from scratch and using the existing weights (warm) of the pre-trained model Mistral 7B v0.1777https://huggingface.co/mistralai/Mistral-7B-v0.1. These four core models were trained on the same amount of data, 64,000 steps of 4 million sub-word tokens each. using identical sets of hyperparameters (see Table 7 in Appendix D). This roughly translates to 3 epochs for the base dataset and 2 for the extended dataset, which according to Muennighoff et al. (2023) is still far from saturating the available data.
@@
Overall, we trained 17 models (7 billion parameters each) using a total of 270,000 GPU-hours. Model training specifications are shown in Table 3. The infrastructure for training included the LUMI supercomputer, Idunn cluster, and Google TPUs through the Tensor Research Cloud program. [^8] To assess the deviation introduced by differences in training infrastructures and platforms across the participating institutions, each team trained a control model with 1.5B parameters based on the Llama 2 architecture. The training setups were identical, utilizing the base dataset. After comparing the validation loss curves from each team, we found that the curves were almost the same, with a deviation of less than $0.05$ in terms of the final convergence validation loss. Besides, we trained two tokenizers with the base and extended datasets separately, both with the same vocabulary size of $32,768$. After an initial test of the fertility [^9] Fertility expresses the fragmentation rate of a tokenizer and is #$\#tokens/\#words$ in one corpus. of the tokenizers, we found the difference between them was only $0.0013$. Therefore, we decided to use the same tokenizer trained with the base dataset for all the models.
@@ 5 Evaluation Framework
    World Knowledge, which assesses the extent of factual information a language model has about the world. This includes historical events, geographical data, scientific facts, cultural knowledge, and more. The model should correctly answer questions or provide information based on real-world knowledge. This skill was evaluated using the NorOpenBookQA [^13] https://huggingface.co/datasets/ltg/noropenbookqa and NRK-Quiz-QA [^14] https://huggingface.co/datasets/ltg/nrk_quiz_qa by Mikhailov et al. (2025).
@@
    Commonsense Reasoning, which involves the ability of a model to make logical inferences based on everyday knowledge and understanding of the world. The model should reason about situations that require practical, everyday knowledge that people take for granted. This skill was evaluated using NorCommonsenseQA [^15] https://huggingface.co/datasets/ltg/norcommonsenseqa Mikhailov et al. (2025) consisting of multiple-choice commonsense question answer-pairs which adapts the corresponding English CommonsenseQA dataset Talmor et al. (2019) to Norwegian.
@@
    Norwegian Language evaluation focuses on the ability of a model to understand and generate text in Norwegian, specifically its grammar, structure, and sentence construction. This skill is important for assessing how well the model handles Norwegian and their specific syntactic rules. It was evaluated using the existing NCB Farsethås and Tjøstheim (2024) and ASK-GEC Jentoft (2023) datasets, and the newly-created NorIdiom [^16] https://huggingface.co/datasets/mimir-project/noridiom dataset.
@@
    Summarization, which measures the ability of a model to condense longer pieces of text into shorter, coherent summaries that capture the main points. This skill is crucial for applications where users need a quick understanding of large volumes of information, such as news articles or research papers. It was evaluated using the NorSumm dataset [^17] https://huggingface.co/datasets/SamiaT/NorSumm Touileb et al. (2025).
@@ 6.1 Core Models
As shown in Table 4 and Figure 2, the performance analysis of core models across various tasks reveals distinct strengths for different configurations. The base (warm-started) configuration consistently excels in Sentiment Analysis, World Knowledge, and Norwegian Language. In contrast, the extended (warm-started) configuration leads in Fairness & Truthfulness, Reading Comprehension, Commonsense Reasoning, Translation, and Variation & Readability, indicating its robust performance for language-intensive tasks. The base configuration generally lags behind others, scoring the lowest across multiple tasks. Meanwhile, the extended configuration performs well, particularly in Summarization. Furthermore, it indicates that we could leverage the existing metadata available at the National Library to tailor subsets of the publisher-controlled copyrighted corpora and build models that excel at specific tasks. However, the difference between the base and extended warm-started models is very small. Further testing is required to assess whether this difference is still statistically significant [^19] Detailed scores available in Appendix F Table 8..
@@
While warm-started models generally outperformed models trained from scratch, there was reduced sensitivity to the presence of copyrighted materials. This suggests that the pre-existing weights, which were primarily trained on English data, diminished the impact of adding high-quality Norwegian copyrighted texts (see more on that in 7).
@@ 7.1 Ethical and Legal Considerations
The use of copyrighted materials in model training raises significant ethical and legal questions. The observed improvements in model quality must be balanced against the rights of content creators, who have not consented to the use of their work. This highlights the need for clear guidelines and fair compensation mechanisms that recognize the value of copyrighted materials in LLM development.
@@ 9 Future Work
Additionally, the observed effects of fiction on model performance highlight the need to 1) examine how different types of fiction –such as fantasy or historical fiction– impact tasks like Sentiment Analysis and Commonsense Reasoning, and 2) design new and adequate benchmarks for evaluating the contribution of fiction in Norwegian LLMs for tasks such as creative writing, plot understanding, or descriptive language use. This targeted investigation could clarify the role of fiction in model training and help refine data curation strategies.
@@ 10 Distribution
The base dataset and models were intended to be freely distributed, as all materials included were granted redistribution permissions under different agreements. After we communicated the results of our investigations to the different partners, some right-holders demanded a reinterpretation of the agreements (primarily the Language Technology Use, Språkteknologiformål), in the light of the results and this new era of generative AI. This prevented us from sharing publicly the exact models trained within the Mímir project, but instead we built a subset of base, which we are calling core, excluding the affected newspapers (around 1B words) and trained models both from scratch and from Mistral 7B v0.1. Their performance is on par with their base counterparts. We are also releasing these models under a permissive license [^20] https://huggingface.co/mimir-project/mimir-mistral-7b-core-scratch and https://huggingface.co/mimir-project/mimir-mistral-7b-core.
@@ Appendix B Sampling
Moreover, from the clean and deduplicated sources, we sub-sampled each non-Norwegian language at an specific sampling ratio until achieving the proportion of documents shown in Figure 5. Pseudo-code for the algorithm used to subsample is shown in Algorithm 1212121https://huggingface.co/mimir-project/mimir-perplexity. We also discovered that a good amount of documents were misclassified by the fastText language identifier Joulin et al. (2016).

### v4

@@ 1 Introduction
However, the vast quantities of data required for training these models often include copyrighted materials, presenting novel challenges related to intellectual property rights and compensation. Additionally, prior research has highlighted significant concerns about dataset composition and quality in large-scale web-crawled datasets, emphasizing the need for more responsible data curation practices (Kreutzer et al., 2022; Artetxe et al., 2022; Penedo et al., 2024). Together, these challenges have led to numerous lawsuits across jurisdictions, fundamentally questioning the legitimacy of training models on copyrighted data without explicit permissions from content creators (Panettieri, 2024; Madigan, 2024; Weisenberger et al., 2024). [^1] See Gervais et al. (2024) for an in-depth introduction on how LLMs are being interpreted in the legal domain.
The first wave of lawsuits emerged shortly after the public release of advanced generative AI models (see Appendix A). Content creators, including authors, visual artists, and musicians, began to express concerns about the unauthorized use of their work in training datasets. Multiple class-action lawsuits were filed in the United States, accusing prominent AI companies such as OpenAI and Meta Platforms of infringing on copyright laws by using copyrighted materials without obtaining explicit permissions. The authors argued that the unauthorized use of their works without any form of compensation or recognition undermines their intellectual property rights and jeopardizes their ability to earn a living from their creative endeavors. In Europe, a coalition of news publishers has taken legal action against Google and Meta Platforms, arguing that the use of journalistic content in training models without fair remuneration constitutes a breach of copyright and undermines the sustainability of high-quality journalism. Likewise, Norwegian rights-holder organizations representing publishing houses across the country, contacted the government in late 2023 expressing their concerns over the use of their material in training generative language models and demanding some sort of compensation were their contents to be used in the training of generative language models. As a result, the Ministry of Culture and Equality (Kultur- og likestillingsdepartemente) instructed the National Library to create a data-driven report they could use in order to make informed decisions in the elaboration of a compensation scheme for the authors. Led by the National Library of Norway, a consortium was formed together with the University of Oslo and the Norwegian University of Science and Technology under the umbrella of the so-called Mímir Project. [^2] A name chosen after a figure in Norse mythology renowned for his knowledge and wisdom.
@@ 3.1 Core Datasets
We followed the recipe of the Norwegian Colossal Corpus (NCC) by Kummervold et al. (2022), adapting and updating it with new up-to-date contents, re-OCRing some materials, enriching their metadata, and ensuring uniform format and functionality across datasets. The preparation involved cleaning, deduplication, metadata tagging, and language balancing to maintain consistent representation of Norwegian, preventing other languages from overshadowing it. The corpus was divided into two main datasets: a base dataset excluding publisher-controlled copyright-protected books and newspapers, [^3] Except for newspapers that fall under the Language Technology Use (Språkteknologiformål), as they were already included in other datasets such as NCC. and an extended dataset that included all collected texts, thus including all of base (see Table 1).
@@
We decided to include texts from other Scandinavian (Swedish, Danish, and Icelandic) and English sources to boost the performance of the resulting language models via cross-lingual transfer (Conneau et al., 2020b; Xue et al., 2021). To ensure that languages other than Norwegian, and primarily coming via Internet crawling, were balanced, we adapted the perplexity-based sampling strategy from De la Rosa et al. (2022) to maintain a high quality in the selected data. Instead of sampling a fixed number of documents, parameters for a Gaussian curve were calculated from 500,000-1M random documents per source, utilizing Wikipedia-based Kneser-Ney language models from Wenzek et al. (2019) and Conneau et al. (2020a). [^4] Built with KenLM Heafield (2011). We also modified the perplexity calculation to account for normalized text. These parameters then guided dataset sub-sampling to target ratios per language, reducing foreign language content while maintaining quality (Appendix B).
@@ 3.2 Domain Specific Subsets
The publisher-controlled copyright-protected materials present in the extended dataset were further divided into groups attending to different criteria. These subsets were carefully designed to test the effect of adding them to the training sets for LLMs. We split the books into fiction vs nonfiction, and original works in Norwegian vs translations. While most books in the collection had metadata information regarding the original language in which a work was written in, genre labels were more scarce. To overcome this limitation, we built a Doc2Vec model Le and Mikolov (2014) that classified fiction vs nonfiction with 98% accuracy and used it to annotate books for which this information was missing. [^5] https://huggingface.co/mimir-project/literary-form-classifier As shown in Table 2, we then built domain specific subsets to investigate 1) the effect of books vs newspapers vs books + newspapers, 2) the effect of factuality by adding only fiction words, only nonfiction works, and nonfiction works + newspapers, and 3) the effect of adding content written originally in Norwegian, such as original books or original books + newspapers, vs translated books.
@@ 3.3 Instruction-tuning Datasets
To align the models more closely with human objectives and assess whether instruction tuning with limited high-quality data can enhance the performance of our pre-trained models across various tasks, we built upon prior work and collected nearly 5,000 instructions annotated by research assistants. [^6] https://huggingface.co/datasets/mimir-project/mimir-instruction The instructions were formatted as (instruction, input, output) triplets, where instruction refers to the directive provided by humans for the model, input is an optional field containing task-related information, and output denotes the desired response that follows the given instruction.
@@ 4 Model Training
    To measure the overall impact of publisher-controlled copyrighted corpora and its effect in realistic scenarios, we conducted pre-training on the base and extended datasets, both from scratch and using the existing weights (warm) of the pre-trained model Mistral 7B v0.1.777https://huggingface.co/mistralai/Mistral-7B-v0.1 These four core models were trained on the same amount of data, 64,000 steps of 4 million sub-word tokens each. using identical sets of hyperparameters (see Table 7 in Appendix D). This roughly translates to 3 epochs for the base dataset and 2 for the extended dataset, which according to Muennighoff et al. (2023) is still far from saturating the available data.
@@
Overall, we trained 17 models (7 billion parameters each) using a total of 270,000 GPU-hours. Model training specifications are shown in Table 3. The infrastructure for training included the LUMI supercomputer, Idunn cluster, and Google TPUs through the Tensor Research Cloud program [^8] To assess the deviation introduced by differences in training infrastructures and platforms across the participating institutions, each team trained a control model with 1.5B parameters based on the Llama 2 architecture. The training setups were identical, utilizing the base dataset. After comparing the validation loss curves from each team, we found that the curves were almost the same, with a deviation of less than $0.05$ in terms of the final convergence validation loss.. Besides, we trained two tokenizers with the base and extended datasets separately, both with the same vocabulary size of $32,768$. After an initial test of the fertility of the tokenizers, [^9] Fertility expresses the fragmentation rate of a tokenizer and is #$\#tokens/\#words$ in one corpus. we found the difference between them was only $0.0013$. Therefore, we decided to use the same tokenizer trained with the base dataset for all the models.
@@ 5 Evaluation Framework
    World Knowledge, which assesses the extent of factual information a language model has about the world. This includes historical events, geographical data, scientific facts, cultural knowledge, and more. The model should correctly answer questions or provide information based on real-world knowledge. This skill was evaluated using the NorOpenBookQA and NRK-Quiz-QA by Mikhailov et al. (2025). [^13] https://huggingface.co/datasets/ltg/noropenbookqa and https://huggingface.co/datasets/ltg/nrk_quiz_qa
@@
    Commonsense Reasoning, which involves the ability of a model to make logical inferences based on everyday knowledge and understanding of the world. The model should reason about situations that require practical, everyday knowledge that people take for granted. This skill was evaluated using NorCommonsenseQA Mikhailov et al. (2025), [^14] https://huggingface.co/datasets/ltg/norcommonsenseqa which consists of multiple-choice commonsense question answer-pairs which adapts the corresponding English CommonsenseQA dataset Talmor et al. (2019) to Norwegian.
@@
    Norwegian Language evaluation focuses on the ability of a model to understand and generate text in Norwegian, specifically its grammar, structure, and sentence construction. This skill is important for assessing how well the model handles Norwegian and their specific syntactic rules. It was evaluated using the existing NCB Farsethås and Tjøstheim (2024) and ASK-GEC Jentoft (2023) datasets, and the newly-created NorIdiom dataset. [^15] https://huggingface.co/datasets/mimir-project/noridiom
@@
    Summarization, which measures the ability of a model to condense longer pieces of text into shorter, coherent summaries that capture the main points. This skill is crucial for applications where users need a quick understanding of large volumes of information, such as news articles or research papers. It was evaluated using the NorSumm dataset Touileb et al. (2025). [^16] https://huggingface.co/datasets/SamiaT/NorSumm
@@ 6.1 Core Models
As shown in Table 4 and Figure 2, the performance analysis of core models across various tasks reveals distinct strengths for different configurations. The base (warm-started) configuration consistently excels in Sentiment Analysis, World Knowledge, and Norwegian Language. In contrast, the extended (warm-started) configuration leads in Fairness & Truthfulness, Reading Comprehension, Commonsense Reasoning, Translation, and Variation & Readability, indicating its robust performance for language-intensive tasks. The base configuration generally lags behind others, scoring the lowest across multiple tasks. Meanwhile, the extended configuration performs well, particularly in Summarization. Furthermore, it indicates that we could leverage the existing metadata available at the National Library to tailor subsets of the publisher-controlled copyrighted corpora and build models that excel at specific tasks. However, the difference between the base and extended warm-started models is very small. Further testing is required to assess whether this difference is still statistically significant. [^18] Detailed scores available in Appendix F Table 8.
@@
While warm-started models generally outperformed models trained from scratch, there was reduced sensitivity to the presence of copyrighted materials. This suggests that the pre-existing weights, which were primarily trained on English data, diminished the impact of adding high-quality Norwegian copyrighted texts (see also Section 7).
@@ 7.1 Ethical and Legal Considerations
The use of copyrighted materials in model training raises significant ethical and legal questions. The observed improvements in model quality must be balanced against the rights of content creators, who have not consented to the use of their work. This highlights the need for guidelines and compensation mechanisms that recognize the value of copyrighted materials in LLM development.
@@ 9 Future Work
Additionally, the observed effects of fiction on model performance highlight the need to 1) examine how different types of fiction –such as fantasy or historical fiction– impact tasks like Sentiment Analysis and Commonsense Reasoning, and 2) design new and adequate benchmarks for evaluating the contribution of fiction in Norwegian LLMs for tasks such as creative writing, plot understanding, or descriptive language use. This investigation could clarify the role of fiction in model training and help refine data curation strategies.
@@ 10 Distribution
The base dataset and models were intended to be freely distributed, as all materials included were granted redistribution permissions under different agreements. After we communicated the results of our investigations to the different partners, some right-holders demanded a reinterpretation of the agreements (primarily the Language Technology Use, Språkteknologiformål), in the light of the results and this new era of generative AI. This prevented us from sharing publicly the exact models trained within the Mímir project, but instead we built a subset of base, which we are calling core, excluding the affected newspapers (around 1B words) and trained models both from scratch and from Mistral 7B v0.1. Their performance is on par with their base counterparts. We are also releasing these models under a permissive license. [^19] https://huggingface.co/mimir-project/mimir-mistral-7b-core-scratch and https://huggingface.co/mimir-project/mimir-mistral-7b-core
@@ Appendix A Legal Cases
  • Silverman v. OpenAI Inc., No. 3:23-cv-03416 (N.D. Cal. July 7, 2023).
  • Thaler v. Perlmutter, 687 F. Supp. 3d 140 (D.D.C. 2023)
  • Doe 1 v. Github, Inc., No. 4:22-cv-06823, 2023 WL 3449131 (N.D. Cal. May 11, 2023)
  • Getty Images (US), Inc. v. Stability AI, Inc., No. 1:23-cv-00135 (D. Del. Feb. 3, 2023)
@@ Appendix B Sampling
Moreover, from the clean and deduplicated sources, we sub-sampled each non-Norwegian language at an specific sampling ratio until achieving the proportion of documents shown in Figure 5. Pseudo-code for the algorithm used to subsample is shown in Algorithm 1.202020https://huggingface.co/mimir-project/mimir-perplexity We also discovered that a good amount of documents were misclassified by the fastText language identifier Joulin et al. (2016).